OpenStudy (anonymous):

log2](4-log2](x) = 3

hartnn (hartnn):

is it \[\log_{2}4-\log_{2}x=3 \]???

OpenStudy (anonymous):

yes

hartnn (hartnn):

u know log a - log b = log (a/b) ??

OpenStudy (anonymous):

yes i think its 32

hartnn (hartnn):

nopes try again....how did u get that?show your work

hartnn (hartnn):

u should get 4/x=8 and not x/4=8